Will someone please help me:
Solve Problems 5, 6, and 7 by completing the square, leaving the results in simplest radical form. Check.
Show Your Work.
Problem #5: x square - 4x = 60
Problem #6: 3x square - 6x = 3
Problem #7: x square + 18x-17 = 0
Thank you very much!

Answer by Alan3354(69443)

Solve Problems 5, 6, and 7 by completing the square, leaving the results in simplest radical form. Check.
Show Your Work.
Problem #5: x square - 4x = 60
x^2 - 4x + 4 = 60+4
(x-2)^2 = 8^2
x-2 = +8, x = 10
------------------
x-2 = -8, x = -6
------------------
Problem #6: 3x square - 6x = 3
Divide by 3, then do it like the first one.
x^2 - 2x = 1
The "c" term is (1/2 the "b" term) squared.
b = 2 --> c = 1
Add 1 to both sides.
x^2 - 2x + 1 = 4 etc
-------------------
Problem #7: x square + 18x-17 = 0
Do it the same way.
Email me at gsihoutx@aol.com if you have trouble, or to check your work.